The pKa of the ammonium ion is 9.25. At what pH is ammonia 35 % ionized?
Answer #1

The fraction of un-ionized ammonia is given by

f = 1/(10^(pKa-pH) + 1)

0.65 = 1/[10^(9.25 - pH) + 1]

  10^(9.25 - pH) + 1 = 1/0.65

pH = 9.52 answer